# supReferences

### v2

@@ 1.2 Contributions of the paper
The second contribution of this paper is to provide a first theoretical analysis of the probability $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$. Considering the same filtering problem as the one described above, we prove that there exists a finite constant $C_{\kappa}$ such that for any $q\in(0,1)$ and time horizon $T\geq 1$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq C_{\kappa}(\log(T)+\log(1/q))$. Consequently, to guarantee that we still have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ as $T$ increases, the number of particles only needs to grow at speed $\log(T)$ while, to decrease $q$, the number of particles should only grow at speed $\log(1/q)$. To the best of our knowledge Del Moral, (2013) provides the only previous attempt to study the ability of PFs to jointly estimate multiple filtering distributions. In particular, in Del Moral, (2013, Corollary 14.3.3, page 435) it is shown that, for some finite constant $C_{\kappa}$, for all $q\in(0,1)$ and $T\in\mathbb{N}$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\geq\kappa)\leq q$ for all $N\geq C_{\kappa}\log(T)\log(1/q)$ and for any function $f$ with oscillations bounded by one. This result is however obtained under strong assumptions on the SSM (which are not verified for the filtering problem that we consider in this work) and it only provides a theoretical guarantee for estimating the filtering expectations $\{\hat{\eta}_{t}(f)\}_{t=1}^{T}$ of a single function.
Despite these reassuring results concerning the time evolution of PF estimates, it is not entirely satisfactory to delegate (life-critical) filtering tasks to algorithms that we know will fail infinitely often. Following a previous discussion, to resolve this limitation of PFs some de-randomization is needed, and a possible approach to do so is to replace its underpinning independent $\mathcal{U}(0,1)$ random variables by randomized quasi-Monte (RQMC) point sets. This idea has been proposed and studied in Gerber and Chopin, (2015) with the aim of obtaining a filtering algorithm, called sequential quasi-Monte Carlo (SQMC), which has a faster convergence rate (as $N\rightarrow\infty$) than PFs. Letting $(\tilde{\eta}^{N}_{t})_{t\geq 1}$ denote the estimate of $(\hat{\eta}_{t})_{t\geq 1}$ obtained with SQMC, the third contribution of this paper is to show, for the one-dimensional filtering problem outlined above, $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\tilde{\eta}^{N}_{t}-\hat{\eta}_{t}\|=0$ with probability one.
@@ 5 Concluding remarks
Given the popularity of PF algorithms in many applications it is important to understand the behaviour of the resulting joint estimation error, $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$. Surprisingly, with the exception of Del Moral, (2013, Corollary 14.3.3, page 435), this latter quantity (or a variant of it) does not seem to have been previously studied in the literature. The results we obtained in this paper are reassuring, but an important question that has yet to be studied is how the dimension of the filtering problem impacts the behaviour of $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.

### v4

@@ 1.2 Contributions of the paper
The second contribution of this paper is to provide a first sharp result on the dependence of the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ to $N$ and $T$. Considering the same filtering problem as the one described above, we prove that there exists a finite constant $C_{\kappa}$ such that for any $q\in(0,1)$ and time horizon $T\geq 1$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq N_{\kappa,T,q}:=C_{\kappa}(\log(T)+\log(1/q))$. For a given $\kappa$, the dependence of $N_{\kappa,T,q}$ to $T$ and $q$ is shown to be sharp, and thus to guarantee that we still have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ as $T$ increases the number of particles only needs to grow at speed $\log(T)$ while, to decrease $q$, the number of particles should only grow at speed $\log(1/q)$. It is worth noting at this stage that Corollary 14.5.7 in Del Moral, (2013) establishes that $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq N^{\prime}_{\kappa,T,q}:=C^{\prime}_{\kappa}(\log(T+e)(1+\log(1/q))$ and for some finite constant $C^{\prime}_{\kappa}$. We also note that it has not been realized (and therefore proved) in this reference that $N$ must increase with $T$ to ensure that $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\geq\kappa)$ remains bounded by $q$ as $T$ increases (and thus that for any fixed $N\geq 1$ we have $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$). We finally remark that this result is obtained under (strong) assumptions which are not verified for the filtering problem that we consider in this work.
Despite the reassuring results we proved regarding the time evolution of PF estimates, it is not entirely satisfactory to delegate (life-critical) filtering tasks to algorithms that we know will fail infinitely often. Following a previous discussion, to resolve this limitation of PFs some de-randomization is needed, and a possible approach to do so is to replace its underpinning independent $\mathcal{U}(0,1)$ random variables by randomized quasi-Monte (RQMC) point sets. This idea has been proposed and studied in Gerber and Chopin, (2015) with the aim of obtaining a filtering algorithm, called sequential quasi-Monte Carlo (SQMC), which has a faster convergence rate (as $N\rightarrow\infty$) than PFs. Letting $(\tilde{\eta}^{N}_{t})_{t\geq 1}$ denote the estimate of $(\hat{\eta}_{t})_{t\geq 1}$ obtained with SQMC, the third contribution of this paper is to show, for the one-dimensional filtering problem outlined above, $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\tilde{\eta}^{N}_{t}-\hat{\eta}_{t}\|=0$ with probability one.
@@ 5 Concluding remarks
Given the popularity of PF algorithms in many applications it is important to understand the behaviour of the resulting joint estimation error, $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$. Surprisingly, with the exception of Del Moral, (2013, Corollary 14.5.7, page 449), this latter quantity (or a variant of it) does not seem to have been previously studied in the literature. The results we obtained in this paper are reassuring, but an important question that has yet to be studied is how the dimension of the filtering problem impacts the behaviour of $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.